A mass of ideal gas of volume 400 cm3 at a temperature of 27 ˚C expands adiabatically until its volume is 500 cm3. Calculate the new temperature. The gas is then compressed isothermally until its pressure returns to the original value. Calculate the final volume of the gas. Assume γ = 1.40.
1
Expert's answer
2020-07-07T10:03:09-0400

"T_1V_1^{\gamma-1}=T_2V_2^{\gamma-1}\to T_2=\frac{T_1V_1^{\gamma-1}}{V_2^{\gamma-1}}=(27+273)(\frac{400\cdot10^{-6}}{500\cdot10^{-6}})^{1.4-1}=274 K"


"p_2V_2=p_3V_3=p_1V_3\to V_3=\frac{p_2V_2}{p_1}"


"p_1V_1^\gamma=p_2V_2^\gamma\to \frac{p_2}{p_1}=\frac{V_1^\gamma}{V_2^\gamma}"


"V_3=\frac{V_1^\gamma}{V_2^\gamma}V_2=(\frac{400\cdot10^{-6}}{500\cdot10^{-6}})^{1.4}\cdot 500\cdot10^{-6}=366\cdot10^{-6}" "m^3=366" "cm^3"